How many nitrogen atoms are present in 2.79 mol of NH3
1
Expert's answer
2020-12-08T05:31:18-0500

Solution:

1 mole of NH3 has 1 mole of N atoms and 3 moles of H atoms.

So, 2.79 mol NH3 have 2.79 mol of N atoms.


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

nitrogen atoms = (2.79 mol of N atoms) × (6.022×1023 atoms / 1 mol) = 16.80×1023 atoms of N

nitrogen atoms = 16.80×1023 atoms or 1.68×1024 atoms


Answer: 1.68×1024 atoms of nitrogen are present in 2.79 mol of NH3.